# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v6

@@ G.3 Proofs of Lemmas for Theorem 3
$\left\{\sum_{i\in S\setminus\{i^{\star}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right\}\gamma\left(v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})\right)\leq\left\{\sum_{i\in S\setminus\{i^{\star}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right\}\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})$
$\Leftrightarrow\left(\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})+\sum_{i\in S\setminus\{i^{\star}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma\right)\left(v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})\right)$
$\leq\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})\left(v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})+\sum_{i\in S\setminus\{i^{\star}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right)$
$\Leftrightarrow\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})+\sum_{i\in S\setminus\{i^{\star}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}.$ (G.3)
Therefore, for all $S\in\mathcal{S}$, we have
$R(S,\mathbf{w}_{V})$ $=\frac{\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})r_{i}}{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}=\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})+\sum_{i\in S\setminus\{i^{\star}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}$
$\leq\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}\leq\frac{\max_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})},$
where the first inequality holds by (G.3), and the last inequality holds since $f(x)=\frac{x}{v_{0}+x}$ is an increasing function.
Case 2. $i^{\star}\notin S$.
Let us return to (G.2). Since $\frac{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\geq 1$ for any $S\in\mathcal{S}$, we have
$\gamma\leq\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\cdot\frac{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})},$
@@
$\frac{\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
Hence, for all $S\in\mathcal{S}$, we get
$R(S,\mathbf{w}_{V})=\frac{\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\max_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
@@ Appendix H Proofs of Theorem 4
We begin the proof by introducing additional useful lemmas. Lemma H.1 shows that $\tilde{R}_{t}(S_{t})$, defined in (6), is an upper bound of the true expected revenue of the optimal assortment, $R_{t}(S_{t}^{\star},\mathbf{w}^{\star})$.
Lemma H.1 (Lemma 4 in Oh and Iyengar 40).
@@ H.1 Proof of Theorem 4
where the inequality holds by the triangle inequality. Now, we bound the terms on the right-hand side of (H.1) individually. For the first term, we have
$\sum_{t=1}^{T}$ $\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|\bar{x}_{ti}-\tilde{x}_{ti}\|_{H_{t}^{-1}}$
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left\|\mathbb{E}_{\mathbf{w}_{t+1}}[x_{tj}]-\mathbb{E}_{\mathbf{w}^{\star}}[x_{tj}]\right\|_{H_{t}^{-1}}$
@@
where the last inequality holds by Lemma E.2. Using similar reasoning, we can bound the second term of (H.1). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
@@
Finally, we bound the third term of (H.1). By the Cauchy-Schwarz inequality, we have
@@
where the last inequality holds by Lemma H.3. Plugging (H.2), (H.3), and (H.4) into (H.1), we get
@@
Now, we bound the term (D). Define $Q:\mathbb{R}^{|S_{t}|}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{|S_{t}|})\in\mathbb{R}^{|S_{t}|}$, $Q(\mathbf{u})=\sum_{i=1}^{|S_{t}|}\frac{\exp(u_{i})}{v_{0}+\sum_{j=1}^{|S_{t}|}\exp(u_{j})}$. Then, we have $\left|\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}\right|\leq\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ since $r_{ti}\in[0,1]$. By following the similar reasoning from the equation (E.3) to (E.5) in Section E.1, we have
@@
where the first inequality holds because $\left|\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}\right|\leq\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$. Combining (H.5) and  (H.6), we derive that
@@
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, since $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K\right)$, we have $\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(Bd\sqrt{T}+\frac{1}{\kappa}d^{2}\right).$
@@ H.2 Proofs of Lemmas for Theorem 4
We prove the result by first showing that for any $i\in S_{t}$, we have $r_{ti}\geq\tilde{R}_{t}(S_{t})$. This can be proven similarly to Lemma G.1. Suppose that there exists $i\in S_{t}$ for which $r_{ti}<\tilde{R}_{t}(S_{t})$. Removing item $i$ from the assortment $S_{t}$ results in a higher expected revenue. Consequently, $S_{t}\neq\mathop{\mathrm{argmax}}_{S\in{\mathcal{S}}}\tilde{R}_{t}(S)$, which contradicts the optimality of $S_{t}$. Thus, we have
@@ I.1 Main Proof of Proposition 1
$\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})$ $=\frac{K\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}{v_{0}+K\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}=\mu(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V}).$
@@
$\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\epsilon}{\sqrt{d}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right).$ (I.2)
@@
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}$ $\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$
@@
For simplicity, let $P=\mathbb{P}_{V}$ and $Q=\mathbb{P}_{V\cup\{j\}}$. Then, we can bound the term $\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|$ in (I.3) for any $V\in\mathcal{V}_{d/4-1}$.
@@
where $\|P-Q\|_{\operatorname{TV}}=\sup_{A}|P(A)-Q(A)|$ | is the total variation distance between $P$ and $Q$, $\operatorname{KL}(P\|Q)=\int(\log\mathrm{d}P/\mathrm{d}Q)\mathrm{d}P$ is s the Kullback-Leibler (KL) divergence between $P$ and $Q$, and the last inequality holds by Pinsker’s inequality. We can derive the instance-dependent bound for the KL divergence term in (I.4) by the following lemma:
@@
Plugging (I.4) into (I.3), we get
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}$ $\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$
@@
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{V}\sum_{i\in S^{\star}}$ $p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$

### v9

@@ G.3 Proofs of Lemmas for Theorem 3
$\left\{\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right\}\frac{\left(v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\right)\nu}{v_{0}+1}\leq\left\{\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right\}\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\nu$
$\Leftrightarrow\left(\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})+\frac{\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+1}\right)\left(v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\right)$
$\quad\,\,\leq\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\left(v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})+\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\right)$
$\Leftrightarrow\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})+\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})/(v_{0}+1)}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})}.$ (G.4)
Therefore, we have
$R(S_{t},\mathbf{w}_{V})$ $=\frac{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})r_{ti}}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}=\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\nu+\sum_{i\in S_{t}\setminus\{i^{\prime}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\cdot\nu/(v_{0}+1)}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}$
$\leq\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})}\leq\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})},$
where the first inequality holds by Equation (G.4).
Case 2. $i^{\prime},i^{\star}\notin S_{t}$.
Let us return to Equation (G.3). Since $\frac{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\geq 1$, we have
$\gamma\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\cdot\frac{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})},$
@@
$\frac{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
Hence, we get
$R(S_{t},\mathbf{w}_{V})$ $=\frac{\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}$
$\leq\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
@@ Appendix H Proofs of Theorem 4
We begin the proof by introducing additional useful lemmas. Lemma H.1 shows that $\tilde{R}_{t}(S_{t})$, defined in Equation (6), is an upper bound of the true expected revenue of the optimal assortment, $R_{t}(S_{t}^{\star},\mathbf{w}^{\star})$.
Lemma H.1 (Lemma 4 in Oh and Iyengar 39).
@@ H.1 Proof of Theorem 4
where the inequality holds by the triangle inequality. Now, we bound the terms on the right-hand side of Equation (H.1) individually. For the first term, we have
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\|\bar{x}_{ti}-\tilde{x}_{ti}\|_{H_{t}^{-1}}=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left\|\mathbb{E}_{\mathbf{w}_{t+1}}[x_{tj}]-\mathbb{E}_{\mathbf{w}^{\star}}[x_{tj}]\right\|_{H_{t}^{-1}}$
@@
where the last inequality holds by Lemma E.2. Using similar reasoning, we can bound the second term of Equation (H.1). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
@@
Finally, we bound the third term of Equation (H.1). By the Cauchy-Schwarz inequality, we have
@@
where the last inequality holds by Lemma H.3. Plugging Equation (H.2), (H.3), and (H.4) into Equation (H.1), we get
@@
Now, we bound the term (D). Define $Q:\mathbb{R}^{|S_{t}|}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{|S_{t}|})\in\mathbb{R}^{|S_{t}|}$, $Q(\mathbf{u})=\sum_{i=1}^{|S_{t}|}\frac{\exp(u_{i})}{v_{0}+\sum_{j=1}^{|S_{t}|}\exp(u_{j})}$. Then, we have $\left|\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}\right|\leq\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ since $r_{ti}\in[0,1]$. By following the similar reasoning from Equation (E.3) to Equation (E.5) in Section E.1, we have
@@
where the first inequality holds because $\left|\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}\right|\leq\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$. Combining Equation (H.5) and (H.6), we derive that
@@
If the boundedness assumption on the parameter is relaxed to $\|\mathbf{w}\|_{2}\leq B$, since $\beta_{t}(\delta)=\mathcal{O}\left(B\sqrt{d}\log t\log K+B^{3/2}\sqrt{d\log K}\right)$ (refer Corollary F.1), we have $\textup{{Reg}}_{T}(\mathbf{w}^{\star})=\tilde{\mathcal{O}}\left(B^{3/2}d\sqrt{T}+\frac{1}{\kappa}B^{3}d^{2}\right).$
@@ H.2 Proofs of Lemmas for Theorem 4
We prove the result by first showing that for any $i\in S_{t}$, we have $r_{ti}\geq\tilde{R}_{t}(S_{t})$. Suppose that there exists $i\in S_{t}$ for which $r_{ti}<\tilde{R}_{t}(S_{t})$. Removing item $i$ from the assortment $S_{t}$ results in a higher expected revenue. Consequently, $S_{t}\neq\mathop{\mathrm{argmax}}_{S\in{\mathcal{S}}}\tilde{R}_{t}(S)$, which contradicts the optimality of $S_{t}$. Thus, we have
@@ I.1 Main Proof of Proposition 1
$\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$ $=\frac{K\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}{v_{0}+K\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}=\mu(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V}).$
@@
$\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]\geq\frac{\kappa^{\star}(\mathbf{w}_{V})}{3}\cdot\frac{\epsilon}{\sqrt{d}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right).$ (I.2)
@@
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}$ $\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
@@
For simplicity, let $P=\mathbb{P}_{V}$ and $Q=\mathbb{P}_{V\cup\{j\}}$. Then, we can bound the term $\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|$ in Equation (I.3) for any $V\in\mathcal{V}_{d/4-1}$.
@@
where $\|P-Q\|_{\operatorname{TV}}=\sup_{A}|P(A)-Q(A)|$ | is the total variation distance between $P$ and $Q$, $\operatorname{KL}(P\|Q)=\int(\log\mathrm{d}P/\mathrm{d}Q)\mathrm{d}P$ is s the Kullback-Leibler (KL) divergence between $P$ and $Q$, and the last inequality holds by Pinsker’s inequality. We can derive the instance-dependent bound for the KL divergence term in Equation (I.4) by the following lemma:
@@
Plugging Equation (I.4) into Equation (I.3), we get
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}$ $\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$
@@
$\frac{1}{\left|\mathcal{V}_{d/4}\right|}\sum_{V\in\mathcal{V}_{d/4}}$ $\mathbb{E}_{V}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\right]$